Prairies Oil Sands Inc. is expected to pay a dividend of $1 in one year. If the dividend growth rate is 2 percent forever and the required return is 10 percent, what should the stock be sold for five years from now?

A. $13.53

B. $13.80

C. $14.08

D. $14.62
Answer #1

This question requires application of constant growth dividend discount model according to which: Po - Divi T-9 Po = Price of

Divo Psr-9

Div1 = $1

Based on basic time value of money,

FV = PV * (1 + r)n

Div6 = Div1 * (1 + 10%)5

Div6 = $1 * (1 + 2%)5 = $1.1041

1.1041 s=n10-0.02

P5 = 13.80(OptionB)